OpenStudy (anonymous):

Simplify square root of 150 a. 5* square root of 6 b. 6* square root of 5 c. 26* square root of 6 d. 36* square root of 5

OpenStudy (lgbasallote):

\[\huge \sqrt{150} \implies \sqrt{25 \times 6}\] do you now how to solve it now?

OpenStudy (anonymous):

not really /:

OpenStudy (anonymous):

@Posideon

OpenStudy (lgbasallote):

\[\huge \sqrt{25 \times 6} \implies \sqrt{25} \times \sqrt 6\] what about now?

OpenStudy (anonymous):

A ?

OpenStudy (ash2326):

@Senior012 in a square root if you have a term like a^2 \[\sqrt {a\times a}=a\] so here we have \[\sqrt{25 \times 6}=\sqrt {5 \times 5\times 6}\] What do you think we'll get now?

OpenStudy (lgbasallote):

are you sure? @Senior012 ? ;)

OpenStudy (anonymous):

yes ? lol

OpenStudy (anonymous):

doood ive solved these like 50 times check out khan academy or something

OpenStudy (lgbasallote):

lol that's right =))